Ch 17 HWw Problem 17.34 Constants I Periodic Table Part A The buner on an electric stove has a power output of 2.0 kW, A 620 g stainless steel tea kettle is filled with 20 c water and placed on the already hot burner If it takes 2.1 min for the water to reach a boil, what volume of water, in cm3,was in the kettle? Stainless steel is mostly iron, so you can assume its specitfic heat is that of iron Express your answer using two significant figures Submit Request Answer Next > Provide Feedback
Answer #1

Solution:

=> 2000 * 2.1 * 60 = m * C_p * dT for iron + m * C_p * dT for water

=> 252000 = 0.62 * 450 * (100-20) + 1000 * volume * 4180 * (100-20)

=> 252000 = 22320 + 334400000*V

=> V = 689 cm^3

please rate thanks.